40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 28 12:56:17 2009, Yoichi, Update, Locking, RF CARM hand off problem 
    Reply  Fri Aug 28 13:49:07 2009, rob, Update, Locking, RF CARM hand off problem 
Message ID: 1960     Entry time: Fri Aug 28 13:49:07 2009     In reply to: 1959
Author: rob 
Type: Update 
Category: Locking 
Subject: RF CARM hand off problem 

Last night, the lock script proceeded to the RF CARM hand-off about half of the time.
However, the hand off was still unsuccessful.

It failed instantly when you turn on the REFL1 input of the CM board, even
when the REFL1 input gain was very low, like -28dB.

I went to the LSC rack and checked the cabling.
The output from the PD11_I (REFL_2) demodulation board is split
into two paths. One goes directly to the ADC and the other one goes
to an SR560. This SR560 is used just as an inverter. Then
the signal goes to the REFL1 input of the CM board.

I found that the SR560 was set to the A-B mode, but B input was open.
This made the signal very noisy. So I changed it to A only mode.
There was also a 1/4 attenuator between the PD11_I output and the SR560.
I took it out and reduced the gain of SR560 from 10 to 2.
These changes allowed me to increase the REFL1 gain to -22dB or so.
But it is still not enough.

I wanted to check the CM open loop TF before the hand-off, but I could
not do that because the lock was lost instantly as soon as I enabled the
test input B of the CM board.
Something is wrong with the board ?

Using the PD11_I signal going into the ADC, I measured the transfer functions
from the CM excitation (digital one) to the REFL_DC (DC CARM signal) and PD11_I.
The TF shapes matched. So the PD11_I signal itself should be fine.

We should try:
* See if flipping the sign of PD11_I signal going to REFL1 input solve the problem.
* Try to measure the CM analog TF again.
* If the noise from the servo analyzer is a problem, try to increase the input gains
of the CM board and reduce the output gain accordingly, so that the signal flowing
inside the CM board is larger.

I'd bet it's in a really twitchy state by the time the script gets to the RF CARM handoff, as the script is not really validated up to that point. It's just the old script with a few haphazard mods, so it needs to be adjusted to accomodate the 15% power drop we've experienced since the last time it was locked.

The CM servo gain needs to be tweaked earlier in the script--you should be able to measure the AO path TF with the arm powers at 30 or so. I was able to do this with the current SR785 setup earlier this week without any trouble.

The 1/4 attenuator is there to prevent saturations on the input to the SR560 when there's still a CARM offset.

Not sure if flipping the sign of PD11 is right, but it's possible we compensated the digital gains and forgot about it. This signal is used for SRCL in the initial acquisition, so we'd have noticed a sign flip.
ELOG V3.1.3-